OpenStudy (anonymous):

Do you need to know Linear Algebra 1 to understand cal 3? I'm trying to start cal 3 ( I'm applying in engineering next semester and I would like to get it out of the way) but I only have cal 1 and cal 2 not linear :S

OpenStudy (anonymous):

In the first lecture video of this course the lecturer does assume the students now some linear algebra when he talks about knowing how to write an equation of a line. \[y = mx + b\] (He doesn't use this equation exactly. I just don't know how to use subscript in the equation editor here. He represents a line's equation as \[y - y0 = m(x - x0)\] where the zeros are subscript.) This happens around the 7 minute mark of the first lecture.